OpenStudy (anonymous):

How do I find the length? A rectangle is drawn so that the width is 2 feet shorter than the length. The area of the rectangle is 63 square feet. Find the length of the rectangle.

OpenStudy (anonymous):

we can calculus the area of rectangle by formula: width*length = area call: w is width and l is length so we have: area = w*l and width is 2 feet shorter than the length so: w = l - 2. sub it to function we have: area = (l-2)*l = 63 so: l^2 - 2l - 63 = 0. From now can you solve for solution?

OpenStudy (anonymous):

@hosiduy so I do 1-21-63?

OpenStudy (anonymous):

Oh no, those are not 1's, they are L's. :D \[L^2-2L-63=0\]

OpenStudy (anonymous):

@agentc0re oops lol I still don't get how do I solve that?

OpenStudy (anonymous):

Quadratic Equation. :D

OpenStudy (anonymous):

So I would add 2L to the other side? @agentc0re

OpenStudy (anonymous):

Oh no no no. The quadratic equation is this: \[\huge x={-b \pm \sqrt{b^2 - 4*a*c} \over 2a}\]

OpenStudy (anonymous):

have you seen that before?

OpenStudy (anonymous):

no ;(

OpenStudy (anonymous):

HELPPPP

OpenStudy (anonymous):

@agentc0re

OpenStudy (anonymous):

well that equation is used when you have an equation in the form: \[\huge ax^2+bx+c\] In this case, you have L instead of x. Make sense?

OpenStudy (anonymous):

a, b and c are the numbers or coefficients in front of the x's, or L's in your case.

OpenStudy (anonymous):

So it's L^2+2L+63 ? @agentc0re

OpenStudy (anonymous):

it is but knowing what i said from above, what is a, b and c?

OpenStudy (anonymous):

Idk Did I do it in the right order? @agentc0re

OpenStudy (anonymous):

I'll give yo a hint... b=2, try to figure out a and c.

OpenStudy (anonymous):

c=63

OpenStudy (anonymous):

IDK what I am doing!

OpenStudy (anonymous):

@agentc0re

OpenStudy (anonymous):

Yup, so b=2, c=63, what about a? I know, thats why i am trying to walk you through this.

OpenStudy (anonymous):

2L !

OpenStudy (anonymous):

So it's 2L+2+63?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

what number is in front of \[\huge L^2\]

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (anonymous):

2L^2+2L+63

OpenStudy (anonymous):

\[\huge L^2+2L+63 =0\] is your answer but you have to solve for L. In order to solve for L you have to use the quadratic equation, which is what i'm trying to teach to you but you're not figuring out how to get a, b and c which is what we need to plug it in to the formula i gave you to get our solution for L. :D It's quite possible that you're solution doesn't need to be L but it's just the polynomial.

OpenStudy (anonymous):

I'm so lost so to find the length of the triangle I have to plug those #'s into the equation...

OpenStudy (anonymous):

This should help. http://www.youtube.com/watch?v=EeVqtpuMFOU

OpenStudy (anonymous):

@agentc0re

OpenStudy (anonymous):

Ok I plugged in the numbers like the guy did in the video but I got -248 you can't sqrt that?!

OpenStudy (anonymous):

Nope, unfortunately you can't. :/ Well you can, but you're going to have part of an answer that is imaginary.

OpenStudy (anonymous):

Good job tho, that's really impressive to have watched that video and then used the quadratic formula. :D

OpenStudy (anonymous):

@agentc0re So how do I find the length of my triangle?

OpenStudy (anonymous):

triangle?? you mean rectangle... why don't you try factoring first..... that would've been faster....:)

OpenStudy (anonymous):

hey desibaybee, did u get this!

OpenStudy (anonymous):

NO :( @cfraser007

OpenStudy (anonymous):

Do you have the answers to this problem? Or is it something you have to enter in online?

OpenStudy (anonymous):

ok...I am trying! lets c how far i get, b4 i start helping u!

OpenStudy (anonymous):

@agentc0re No thats exactly how it is and I have to fill in the blank!

OpenStudy (anonymous):

Then the answer you got by using the quadratic equation is correct. Even though it's not a real length, it is the correct answer unless there's been a typo of the problem along the way. There is no way to factor this so that it's solvable (at least I don't see it)

OpenStudy (anonymous):

I don't think you are suppose to use that formula for this because I didn't learn that in class? @agentc0re

OpenStudy (anonymous):

What level of Math class is this?

OpenStudy (anonymous):

what did u learn...or what methods? or what was the lesson you all were talking about before seeing this problem!

OpenStudy (anonymous):

algebra but like development 2

OpenStudy (anonymous):

factoring stuff like that but I remember him just making an equation out of it

OpenStudy (anonymous):

haha! this problem is so simple guys! v r bunch of too smart folks trying to overthink it! lol...ok solution here it is! area of a rectangle is 63...and the width is 2 feet shorter than the length! so...the difference between 2 numbers must b 2 and multiply to equal 63! the only set is 7 and 9! since 7 x 9 = 63 and 9 - 7 = 2! I think, now u can figure out which one is width and which one is length! Enjoy!

OpenStudy (anonymous):

@cfraser007 yeah when I factored it I got 7 and 9 but witch is the length and which is the width?

OpenStudy (anonymous):

So is the width 7?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

@Desibaybee we had a sign messed up. the equation needed to be: \[\huge x^2-2x-63\] solve that using the equation now, you'll see that you get those answers.

OpenStudy (anonymous):

@cfraser007 WOOOOO THANK YOU SOOO MUCH OMGSHHH THAT TOOK ME FOREVER SO SIMPLE!!! ALL WE HAD TO DO WAS FACTOR THE EQUATION!!! :D

OpenStudy (anonymous):

i'm sorry i didn't catch the sign error before. :/

OpenStudy (anonymous):

@agentc0re It's ok hun thanks so much for trying and helping tho! :)

OpenStudy (anonymous):

NO NO NO, this is not factoring!!! i promise you this is not factoring.

OpenStudy (anonymous):

Well when I factored it and put 9 for the length it was right! @agentc0re

OpenStudy (anonymous):

factoring is when you have\[\huge x^2+2x+1 = (x+1)(x+1)\] Foiling, is when you do the reverse. I'm just trying to help with the terminology. Anyways, glad you got the answer. :D

OpenStudy (anonymous):

agentcOre, I see where u r getting at! u get (L+7)(L-9) = 0, then L= -7 and L= 9! May b u r concerned about the negative signs! but it should not b a problem...just say that, v take absolute values since v r dealing with lengths and distances... ! problem solved! (i think!)

OpenStudy (anonymous):

oh! my bad! agentcOre!

OpenStudy (anonymous):

@cfraser007 Ya, that is factoring but i didn't see that being done until just now. There was a math error on my part when trying to explain the quadratic formula and i dropped a minus sign when i shouldn't have. I think we're all confused. hahahhaha. OH well.

OpenStudy (anonymous):

haha! lol